an isotope of Strontium 90 has a half life of 38 years. if you have 500mg of this isotope today how many milligrams will you have in 19 years
Answer by ankor@dixie-net.com(22740)

an isotope of Strontium 90 has a half life of 38 years.
if you have 500mg of this isotope today how many milligrams will you have in 19 years.
:
The radioactive decay formula:
A = Ao*2(-t/h), where:
A = amt after t yrs
Ao = initial amt
t = time
h = half-life of substance
:
A = 500*2(-19/38)
A = 500*2(-1/2}
A = 500 * .7071
A = 353.55 milligrams after 19 yrs
: